# Diffusion Transformer Captures Spatial-Temporal Dependencies: A Theory for Gaussian Process Data

### v1

@@ 1 Introduction
Recently, the astounding success in high-fidelity video generation strengthens the seemingly unlimited potentiality of diffusion models (Gupta et al., 2023; Liu et al., 2024b). These models advocate transformers over the traditional U-Net for parameterizing the score function. A high-level intuition is that video data comprises rich spatial and temporal dependencies, induced by the underlying dynamics of objects and background. For example, the movement of an object should be continuous along the time. These dependencies resonate well with the self-attention mechanism in transformers for capturing token-wise correlation, suggesting benefits for learning with sequential video data. We illustrate diffusion transformer accurately learning spatial-temporal dependencies in Figure 1.
Figure 1: Diffusion transformer learns spatial-temporal dependencies. The diffusion transformer is trained with data sampled from a stationary Gaussian process consisting of $128$ time steps. At each time step, the data dimension is $8$. We obtain 1000 generated samples at each time step. The left large heat map demonstrates the estimated temporal correlation (see Appendix D for the estimation method) in the process between different time steps, which aligns well with the ground truth on the right. The smaller heat maps are the estimated covariance matrix of data at a single time step, which demonstrate the spatial dependencies in data. They also align well with the ground truth.
Despite the empirical success, there lacks a rigorous understanding of diffusion transformers for sequential data modeling. Different from static data, such as images, sequential data consists of a series of dependent data frames. The data sequence can be extensively long. For instance, a one-minute video would contain over 1440 image frames. Therefore, naïvely handling the sequential data by stacking all the frames as if independent will lead to a large dimensionality dependence and the spatial-temporal correlation further complicates matters. As a result, existing results of diffusion models for static data can provide highly limited insights on the following fundamental questions:
@@
We answer these questions firmly by studying for the first time of using diffusion transformers for learning Gaussian process data. Besides its simplicity, Gaussian process exhibits intriguing and salient properties. Firstly, Gaussian process data can be high-dimensional, highlighting the influence of data dimensionality in diffusion transformer. Secondly, the spatial-temporal dependencies are the defining quantities of a Gaussian process. This necessitates an effective learning of these dependencies. In fact, Gaussian process can encode a wide variety of complicated correlations in real-world applications (Seeger, 2004; Williams and Rasmussen, 2006). For instance, Brownian motion falls into the category of Gaussian process for describing particle movements in a fluid. Gaussian process is also a powerful statistical tool for regression, classification and forecasting problems (Banerjee et al., 2013; Deringer et al., 2021; Chen et al., 2021; Borovitskiy et al., 2021).
Contributions  Our results show that by construction, transformers can adapt to the spatial-temporal dependencies so as to promote the learning efficiency. Furthermore, we show sample complexity bounds of diffusion transformers, demonstrating the influence of the decay of correlation in the sequential data. We summarize our contributions as follows.
@@
$\bullet$ Built upon our score function approximation theory, we establish the first sample complexity bound for diffusion transformers in learning sequential data (Theorem 2). We show that the generalization error scales with $1/\sqrt{n}$, where $n$ is the sample size. It also depends on the decay of temporal dependencies and long-horizon dependencies introduces large dependency on the sequence length.
@@
On the empirical side, Diffusion Transformer (DiT) (Peebles and Xie, 2023) challenges the common choice of U-Net for image generation, providing state-of-the-art performance. Moreover, diffusion transformer exhibits appealing scalability towards better generation qualities with larger model sizes. More recently, diffusion transformers are leveraged in video generation (Gupta et al., 2023; Liu et al., 2024b), where the video data is patchified and rearranged into a long sequence. Besides, transformers are also used in discrete diffusion models (Sun et al., 2022; Austin et al., 2021).
@@
We use bold letters to denote vectors and matrices. For a vector $\mathbf{v}$, we denote its Euclidean norm as $\left\lVert\mathbf{v}\right\rVert_{2}$. For a matrix $\mathbf{A}$, we denote its operator, Frobenius norm as $\left\lVert\mathbf{A}\right\rVert_{2}$ and $\left\lVert\mathbf{A}\right\rVert_{\rm F}$, respectively. Moreover, we denote $\left\lVert\mathbf{A}\right\rVert_{\infty}=\max_{i,j}{\left|\mathbf{A}\right|}$. We denote the condition number of a positive definite matrix $\mathbf{A}$ by $\kappa(\mathbf{A})=\lambda_{\max}(\mathbf{A})/\lambda_{\min}(\mathbf{A})$, where $\lambda_{\max}$ and $\lambda_{\min}$ denote the maximum and minimum eigenvalues. We denote $f\lesssim g$ when there exists a constant $C>0$ with $f\leq Cg$.
@@ 2 Gaussian Process and Diffusion Transformer
We denote $\{\mathbf{X}_{h}\}_{h\in[0,H]}$ as a continuous-time Gaussian process in the time interval $[0,H]$. The process lives in the $d$-dimensional Euclidean space, i.e., $\mathbf{X}_{h}\in\mathbb{R}^{d}$ for any $h\in[0,H]$. A defining property of Gaussian process is that for any finite collection of time indexes $h_{1},\dots,h_{N}$ with $N\in\mathbb{N}^{+}$, the joint distribution of $\{\mathbf{X}_{h_{1}},\dots,\mathbf{X}_{h_{N}}\}$ is still Gaussian. As a particular example, for a fixed time $h$, the marginal distribution of $\mathbf{X}_{h}$ is Gaussian.
To fully describe a Gaussian process, we need the concept of mean and covariance functions. Roughly speaking, a mean function $\bm{\mu}(h)=\mathbb{E}[\mathbf{X}_{h}]$ characterizes the expected evolution trend of the process. A covariance function $\bm{\Gamma}(h_{1},h_{2})=\mathbb{E}[(\mathbf{X}_{h_{1}}-\bm{\mu}(h_{1}))(\mathbf{X}_{h_{2}}-\bm{\mu}(h_{2}))^{\top}]$ captures the correlation between two time indexes in the process. Note that, when $h_{1}=h_{2}=h$, the covariance function computes the covariance matrix of $\mathbf{X}_{h}$. Remarkably, the covariance function determines many basic properties of the continuous-time process, such as its stationarity, periodicity, and smoothness (Williams and Rasmussen, 2006). In our later study, we reveal an intimate connection between the behavior of the covariance function to the learning efficiency of diffusion transformer.
Throughout the paper, we focus on (strictly) stationary Gaussian processes, that is, the covariance function only depends on the gap between time indexes. Accordingly, we reparameterize the covariance function as $\bm{\Gamma}(h_{1},h_{2})=\gamma(h_{1},h_{2})\bm{\Sigma}$, where $\gamma(\cdot,\cdot)$ is a scalar-output function and $\bm{\Sigma}=\operatorname{\rm Cov}[\mathbf{X}_{h}]$ (identical for any $h$).
@@
In real-world scenarios, an underlying continuous-time process is often perceived by a sequence of data sampled at discrete times. For example, a video typically consists of 24 to 30 image frames per second. When played back, these frames appear seamless to the human eye, which cannot distinguish between individual frames as if the video is continuous. Following the same spirit, we denote $h_{1},\dots,h_{N}$ for a sufficiently large $N\in\mathbb{N}^{+}$ as a uniform grid on the time interval $[0,H]$ and form a discrete sequence $\{\mathbf{X}_{h_{1}},\dots,\mathbf{X}_{h_{N}}\}$ observed at those times from an underlying Gaussian process. By the definition of Gaussian process, if we stack $\mathbf{X}_{h_{1}},\dots,\mathbf{X}_{h_{N}}$ consecutively as a vector in $\mathbb{R}^{dN}$, it follows a Gaussian distribution. The mean is $\bm{\mu}=[\bm{\mu}_{1}^{\top},\dots,\bm{\mu}_{N}^{\top}]^{\top}$ with $\bm{\mu}_{i}=\bm{\mu}(h_{i})$ and the covariance is a block-wise matrix represented as $\bm{\Gamma}\otimes\bm{\Sigma}$, where
@@
Here, $\otimes$ is the matrix Kronecker product. Notably, $\bm{\Gamma}$ captures the temporal dependency across time and $\bm{\Sigma}$ captures the spatial dependency of entries in $\mathbf{X}_{h}$. Since $h_{1},\dots,h_{N}$ form a uniform grid, $\bm{\Gamma}$ is a symmetric Toeplitz matrix with entries taking at most $N$ different values.
@@
Here, $\mathbf{W}_{t}$ and $\overline{\mathbf{W}}_{t}$ are independent Wiener processes and $T$ is a finite terminal time. The initial distribution of $\mathbf{X}_{0}$ is $P_{0}$, which is also the clean data distribution, and we denote $P_{t}$ as the marginal distribution of $\mathbf{X}_{t}$. Since we are corrupting $P_{0}$ by Gaussian noise, $P_{t}$ has a density function $p_{t}$. Thus, $\nabla\log p_{t}$ in the backward process is recognized as the score function, which is unknown and requires learning. Typically, the score function will be parameterized by a neural network and trained by optimizing a loss function, which we will introduce in Section 5. When generating new samples, we simulate a discretized version of the backward process using the learned score function.
@@
Here, $\mathbf{W}_{1},\mathbf{W}_{2}$ are weight matrices, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are offset vectors, $\mathbf{1}$ denotes a vector of ones, and the ReLU activation function is applied entrywise. For our study, the raw input to a transformer is $N$ patches of $d$-dimensional vectors and time $t$ in the backward process. We refer to ${\mathcal{T}}(D,L,M,B,R)$ as a transformer architecture defined by
@@
See Figure 2 for an illustration and explanation of $f_{\rm in}$ and $f_{\rm out}$. For attention layers, we consider ReLU activation for technical convenience and postpone a discussion with softmax activation to Appendix B.5.
Figure 2: Transformer architecture. Here $f_{\rm in}$ is a linear layer to lift input patch to $\mathbb{R}^{D}$, which appends the input raw data with time embedding and other useful information. After passing through $L$ transformer blocks, $f_{\rm out}$ projects each patch into the data original dimension $\mathbb{R}^{d}$ and clip the output range by $R$. We allow the output range to be time $t$ dependent (denoted as $R_{t}$).
@@ 3.1 Score Function for Gaussian Process Data
Recall from Section 2 that the joint distribution of our $N$-patch Gaussian process data is still Gaussian. We can show that at time $t\in(0,T]$, given $\mathbf{v}_{t}=[\mathbf{x}_{1,t}^{\top},\dots,\mathbf{x}_{N,t}^{\top}]^{\top}$ in the backward process, the score function is
$\nabla\log p_{t}(\mathbf{v}_{t})=(\alpha_{t}^{2}\bm{\Gamma}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})\quad\text{for}\quad\alpha_{t}=e^{-t/2}~{}\text{and}~{}\sigma_{t}=\sqrt{1-e^{-t}},$ (2)
where $\bm{\mu}$ is the mean vector. We defer the derivation to Appendix A.1. Observe that the score function for the $i$-th patch depends on the evolution of all patches, reflecting temporal dependencies in sequential data. Moreover, the influence of each patch is determined by the covariance function $\bm{\Gamma}_{ij}=\gamma(h_{i},h_{j})$. In the extreme case of $\bm{\Gamma}_{ij}=\mathds{1}\{i=j\}$, i.e., there is no correlation between patches, the score function reduces to $(\alpha_{t}^{2}\mathbf{I}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$, which isolates patches and reproduces the score function of the Gaussian distribution ${\sf N}(\bm{\mu}_{i},\bm{\Sigma})$ for each patch.
@@ 3.2 Score Function Is the Optimizer of a Convex Function and Gradient Descent Finds It
We consider a fixed time $t\in(0,T]$. It is straightforward to check by the first-order optimality that $\nabla\log p_{t}(\mathbf{v}_{t})$ is the minimizer of the following quadratic objective function,
$\nabla\log p_{t}(\mathbf{v}_{t})=\mathop{\mathrm{argmin}}_{\mathbf{s}\in\mathbb{R}^{dN}}~{}\mathcal{L}_{t}(\mathbf{s}):=\frac{1}{2}\mathbf{s}^{\top}\left(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)\mathbf{s}-(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})^{\top}\mathbf{s}.$ (3)
@@
$\mathbf{s}^{(k+1)}=\mathbf{s}^{(k)}-\eta_{t}\cdot\left(\left(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)\mathbf{s}^{(k)}-(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})\right),$ (4)
@@
For an arbitrarily fixed time $t\in[0,T]$ and $\mathbf{v}_{t}$, given an error tolerance $\epsilon>0$ and any integer $J<N$, if $\bar{\bm{\Gamma}}$ with $\bar{\bm{\Gamma}}_{ij}=\bm{\Gamma}_{ij}\mathds{1}\{|i-j|<J\}$ is positive semidefinite, then running gradient descent in (4) with a suitable step size $\eta_{t}$ for $K=\mathcal{O}{\left(\kappa_{t}\log{\left(1/{\epsilon}\right)}\right)}$ iterations gives rise to

### v2

@@ 1 Introduction
Recently, the astounding success of applying diffusion models in dynamic (sequential) data, including video generation (Gupta et al., 2023; Liu et al., 2024b) and financial data augmentation (Gao et al., 2024), strengthens the seemingly unlimited potentiality of diffusion models . These models advocate transformers over the traditional U-Net for parameterizing the score function. A high-level intuition is that video data comprises rich spatial and temporal dependencies, induced by the underlying dynamics of objects and background. For example, the movement of an object should be continuous along the time. These dependencies resonate well with the self-attention mechanism in transformers for capturing token-wise correlation, suggesting benefits for learning with sequential data. We illustrate diffusion transformer accurately learning spatial-temporal dependencies in Figure 1.
Figure 1: Diffusion transformer learns spatial-temporal dependencies. The diffusion transformer is trained with data sampled from a stationary Gaussian process consisting of $128$ time steps. At each time step, the data dimension is $8$. We obtain 1000 generated samples at each time step. The left large heat map demonstrates the estimated temporal correlation (see Appendix D.1 for the estimation method) in the process between different time steps, which aligns well with the ground truth on the right. The smaller heat maps are the estimated covariance matrix of data at a single time step, which demonstrate the spatial dependencies in data. They also align well with the ground truth.
Despite the empirical success, there lacks a rigorous understanding of diffusion transformers for sequential data modeling. Different from static data, sequential data consists of a series of dependent data frames. The data sequence can be extensively long. For instance, a one-minute video would contain over 1440 image frames, and intraday data in financial applications may be even longer.
Therefore, naïvely treating the entire sequential data solely as high-dimensional data without considering its inner spatial-temporal correlations will lead to a large dimensionality dependence and inefficient learning. As a result, existing results of diffusion models for static data can provide few insights on the following fundamental questions:
@@
We answer these questions for the first time by studying using diffusion transformers for learning Gaussian process data. Besides its simplicity, Gaussian process exhibits intriguing and salient properties. Firstly, Gaussian process data can be high-dimensional, highlighting the influence of data dimensionality in diffusion transformer. Secondly, the spatial-temporal dependencies are the defining quantities of a Gaussian process. This necessitates an effective learning of these dependencies. In fact, Gaussian process can encode a wide variety of complicated correlations in real-world applications (Seeger, 2004; Williams and Rasmussen, 2006). For instance, Brownian motion falls into the category of Gaussian process for describing particle movements in a fluid. Gaussian process is also a powerful statistical tool for regression, classification and forecasting problems (Banerjee et al., 2013; Deringer et al., 2021; Chen et al., 2021; Borovitskiy et al., 2021).
Contributions
Our results show that by construction, transformers can adapt to the spatial-temporal dependencies so as to promote the learning efficiency. Furthermore, we show sample complexity bounds of diffusion transformers, demonstrating the influence of the decay of correlation in the sequential data. We summarize our contributions as follows.
@@
$\bullet$ Built upon our score function approximation theory, we establish the first sample complexity bound for diffusion transformers in learning sequential data (Theorem 2). We show that the generalization error scales with $1/\sqrt{n}$, where $n$ is the sample size. Our generalization error also demonstrates the influence of dependency decay speed and the length of sequences.
@@
On the empirical side, Diffusion Transformer (DiT) (Peebles and Xie, 2023) challenges the common choice of U-Net for image generation, providing state-of-the-art performance. Moreover, diffusion transformer exhibits appealing scalability towards better generation qualities with larger model sizes. More recently, diffusion transformers are leveraged in video generation (Gupta et al., 2023; Liu et al., 2024b), where the video data is patchified and rearranged into a long sequence. Besides, diffusion transformers are also used for other sequential data (Sun et al., 2022; Austin et al., 2021; Campbell et al., 2022; Gao et al., 2024), such as language, music and financial data.
@@
We use bold letters to denote vectors and matrices. For a vector $\mathbf{v}$, we denote its Euclidean norm as $\left\lVert\mathbf{v}\right\rVert_{2}$. For a matrix $\mathbf{A}$, we denote its operator, Frobenius norm as $\left\lVert\mathbf{A}\right\rVert_{2}$ and $\left\lVert\mathbf{A}\right\rVert_{\rm F}$, respectively. Moreover, we denote $\left\lVert\mathbf{A}\right\rVert_{\infty}=\max_{i,j}{\left|\mathbf{A}_{i,j}\right|}$. We denote the condition number of a positive definite matrix $\mathbf{A}$ by $\kappa(\mathbf{A})=\lambda_{\max}(\mathbf{A})/\lambda_{\min}(\mathbf{A})$, where $\lambda_{\max}$ and $\lambda_{\min}$ denote the maximum and minimum eigenvalues. We denote $f\lesssim g$ if $f\leq Cg$ holds for a constant $C>0$.
@@ 2 Gaussian Process and Diffusion Transformer
We denote $\{\mathbf{X}_{h}\}_{h\in[0,H]}$ as a continuous-time Gaussian process in the time interval $[0,H]$. The process lives in the $d$-dimensional Euclidean space, i.e., $\mathbf{X}_{h}\in\mathbb{R}^{d}$ for any $h\in[0,H]$. A defining property of Gaussian process is that for any finite collection of time indices $h_{1},\dots,h_{N}$ with $N\in\mathbb{N}^{+}$, the joint distribution of $\{\mathbf{X}_{h_{1}},\dots,\mathbf{X}_{h_{N}}\}$ is still Gaussian. As a particular example, for a fixed time $h$, the marginal distribution of $\mathbf{X}_{h}$ is Gaussian.
To fully describe a Gaussian process, we need the concept of mean and covariance functions. Roughly speaking, a mean function $\bm{\mu}(h)=\mathbb{E}[\mathbf{X}_{h}]$ characterizes the expected evolution trend of the process. A covariance function $\bm{\Gamma}(h_{1},h_{2})=\mathbb{E}[(\mathbf{X}_{h_{1}}-\bm{\mu}(h_{1}))(\mathbf{X}_{h_{2}}-\bm{\mu}(h_{2}))^{\top}]$ captures the correlation between two time indices in the process. Note that, when $h_{1}=h_{2}=h$, the covariance function computes the covariance matrix of $\mathbf{X}_{h}$. Remarkably, the covariance function determines many basic properties of the continuous-time process, such as its stationarity, periodicity, and smoothness (Williams and Rasmussen, 2006). In our later study, we reveal an intimate connection between the behavior of the covariance function to the learning efficiency of diffusion transformer.
Throughout the paper, we focus on Gaussian processes whose covariance functions only depend on the gap between time indices. Accordingly, we reparameterize the covariance function as $\bm{\Gamma}(h_{1},h_{2})=\gamma(h_{1},h_{2})\bm{\Sigma}$, where $\gamma(\cdot,\cdot)$ is a scalar-output function and $\bm{\Sigma}=\operatorname{\rm Cov}[\mathbf{X}_{h}]$ (identical for any $h$).
@@
In real-world scenarios, an underlying continuous-time process is often perceived by a sequence of data sampled at discrete times. For example, a video typically consists of 24 to 30 image frames per second. When played back, these frames appear seamless to the human eye, which cannot distinguish between individual frames as if the video is continuous. Following the same spirit, we denote $h_{1},\dots,h_{N}$ for a sufficiently large $N\in\mathbb{N}^{+}$ as a uniform grid on the interval $[0,H]$ and form a discrete sequence $\{\mathbf{X}_{h_{1}},\dots,\mathbf{X}_{h_{N}}\}$ observed at those time indices from an underlying Gaussian process. By the definition of Gaussian process, if we stack $\mathbf{X}_{h_{1}},\dots,\mathbf{X}_{h_{N}}$ consecutively as a vector in $\mathbb{R}^{dN}$, it follows a Gaussian distribution. The mean is $\bm{\mu}=[\bm{\mu}_{1}^{\top},\dots,\bm{\mu}_{N}^{\top}]^{\top}$ with $\bm{\mu}_{i}=\bm{\mu}(h_{i})$ and the covariance is a block-wise matrix represented as $\bm{\Gamma}\otimes\bm{\Sigma}$, where
@@
Here, $\otimes$ is the matrix Kronecker product. Notably, $\bm{\Gamma}$ captures the temporal dependency between time indices and $\bm{\Sigma}$ captures the spatial dependency of entries in $\mathbf{X}_{h}$. Since $h_{1},\dots,h_{N}$ form a uniform grid, $\bm{\Gamma}$ is a symmetric Toeplitz matrix with entries taking at most $N$ different values.
Gaussian process data provides explicit description of the spatial-temporal dependencies, but still remains highly relevant to real-world diffusion models. These models utilize a pre-trained Variational AutoEncoder (VAE) to map data into a low-dimensional latent representation (Wang et al., 2023; Blattmann et al., 2023). The typical prior distribution of the low-dimensional representation in VAEs is assumed to be Gaussian. Empirical results have demonstrated the effectiveness of Gaussian latent prior for sequential data modeling in some variants of VAEs (Casale et al., 2018; Fortuin et al., 2020). Our study aligns well with these empirical observations.
@@
Here, $\mathbf{W}_{t}$ and $\overline{\mathbf{W}}_{t}$ are independent Wiener processes and $T$ is a finite terminal diffusion timestep. The initial distribution of $\mathbf{X}_{0}$ is $P_{0}$, which is also the clean data distribution, and we denote $P_{t}$ as the marginal distribution of $\mathbf{X}_{t}$. Since we are corrupting $P_{0}$ by Gaussian noise, $P_{t}$ has a density function $p_{t}$. Thus, $\nabla\log p_{t}$ in the backward process is recognized as the score function, which is unknown and requires learning. Typically, the score function will be parameterized by a neural network and trained by optimizing a loss function, which we will introduce in Section 5. When generating new samples, we simulate a discretized version of the backward process using the learned score function.
@@
Here, $\mathbf{W}_{1},\mathbf{W}_{2}$ are weight matrices, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are offset vectors, $\mathbf{1}$ denotes a vector of ones, and the ReLU activation function is applied entrywise. For our study, the raw input to a transformer is $N$ patches of $d$-dimensional vectors and diffusion timestep $t$ in the backward process. We refer to ${\mathcal{T}}(D,L,M,B,R)$ as a transformer architecture defined by
@@
See Figure 2 for an illustration of $f_{\rm in}$ and $f_{\rm out}$. For attention layers, we consider ReLU activation for technical convenience and postpone a discussion with softmax activation to Appendix B.5.
Figure 2: Transformer architecture. Here $f_{\rm in}$ is a linear layer to lift input patch to $\mathbb{R}^{D}$, which appends the input raw data with time index embedding and other useful information. After passing through $L$ transformer blocks, $f_{\rm out}$ projects each patch into the data original dimension $\mathbb{R}^{d}$ and clip the output range by $R$. We allow the output range to be diffusion timestep $t$ dependent (denoted as $R_{t}$).
@@ 3.1 Score Function for Gaussian Process Data
Recall from Section 2 that the joint distribution of our $N$-patch Gaussian process data is still Gaussian. We can show that at diffusion timestep $t\in(0,T]$, given $\mathbf{v}_{t}=[\mathbf{x}_{1,t}^{\top},\dots,\mathbf{x}_{N,t}^{\top}]^{\top}$ in the backward process, the score function is
$\nabla\log p_{t}(\mathbf{v}_{t})=-(\alpha_{t}^{2}\bm{\Gamma}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})\quad\text{for}\quad\alpha_{t}=e^{-t/2}~{}\text{and}~{}\sigma_{t}=\sqrt{1-e^{-t}},$ (2)
where $\bm{\mu}$ is the mean vector, and $\alpha_{t}$ and $\sigma_{t}$ are determined by the diffusion forward process. We defer the derivation to Appendix A.1. Observe that the score function for the $i$-th patch depends on the evolution of all patches, reflecting temporal dependencies in sequential data. Moreover, the influence of each patch is determined by the covariance function $\bm{\Gamma}_{ij}=\gamma(h_{i},h_{j})$. In the extreme case of $\bm{\Gamma}_{ij}=\mathds{1}\{i=j\}$, i.e., there is no correlation between patches, the score function reduces to $(\alpha_{t}^{2}\mathbf{I}\otimes\bm{\Sigma}+\sigma_{t}^{2}\mathbf{I})^{-1}(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})$, which isolates patches and reproduces the score function of the Gaussian distribution ${\sf N}(\bm{\mu}_{i},\bm{\Sigma})$ for each patch.
@@ 3.2 Score Function Is The Optimizer of A Convex Function And Gradient Descent Finds It
We consider a fixed diffusion timestep $t\in(0,T]$. It is straightforward to check by the first-order optimality that $\nabla\log p_{t}(\mathbf{v}_{t})$ is the minimizer of the following quadratic objective function,
$\nabla\log p_{t}(\mathbf{v}_{t})=\mathop{\mathrm{argmin}}_{\mathbf{s}\in\mathbb{R}^{dN}}~{}\mathcal{L}_{t}(\mathbf{s}):=\frac{1}{2}\mathbf{s}^{\top}\left(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)\mathbf{s}+(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})^{\top}\mathbf{s}.$ (3)
@@
$\mathbf{s}^{(k+1)}=\mathbf{s}^{(k)}-\eta_{t}\cdot\left(\left(\alpha_{t}^{2}(\bm{\Gamma}\otimes\bm{\Sigma})+\sigma_{t}^{2}\mathbf{I}\right)\mathbf{s}^{(k)}+(\mathbf{v}_{t}-\alpha_{t}\bm{\mu})\right),$ (4)
@@
For an arbitrarily fixed $t\in[0,T]$ and $\mathbf{v}_{t}$, given an error tolerance $\epsilon>0$ and any integer $J<N$, if $\bar{\bm{\Gamma}}$ with $\bar{\bm{\Gamma}}_{ij}=\bm{\Gamma}_{ij}\mathds{1}\{|i-j|<J\}$ is positive semidefinite, then running gradient descent in (4) with a suitable step size $\eta_{t}$ for $K=\mathcal{O}{\left(\kappa_{t}\log{\left(1/{\epsilon}\right)}\right)}$ iterations gives rise to